### v2

# From Lab to Field: Real-World Evaluation of an AI-Driven Smart Video Solution to Enhance Community Safety

Abstract
This article adopts and evaluates an AI-enabled Smart Video Solution (SVS) designed to enhance safety in the real world. The system integrates with existing infrastructure camera networks, leveraging recent advancements in AI for easy adoption. Prioritizing privacy and ethical standards, pose based data is used for downstream AI tasks such as anomaly detection. Cloud-based infrastructure and mobile app are deployed, enabling real-time alerts within communities. The SVS employs innovative data representation and visualization techniques, such as the Occupancy Indicator, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to understand pedestrian behaviors and enhance public safety. Evaluation of the SVS demonstrates its capacity to convert complex computer vision outputs into actionable insights for stakeholders, community partners, law enforcement, urban planners, and social scientists. This article presents a comprehensive real-world deployment and evaluation of the SVS, implemented in a community college environment across 16 cameras. The system integrates AI-driven visual processing, supported by statistical analysis, database management, cloud communication, and user notifications. Additionally, the article evaluates the end-to-end latency from the moment an AI algorithm detects anomalous behavior in real-time at the camera level to the time stakeholders receive a notification. The results demonstrate the system’s robustness, effectively managing 16 CCTV cameras with a consistent throughput of 16.5 frames per second (FPS) over a 21-hour period and an average end-to-end latency of 26.76 seconds between anomaly detection and alert issuance.
Introduction
Figure 1: The end-to-end system architecture coupled with notification’s data flow. $S_{1}$ represents the first scene that detects a behavior anomaly. The yellow line shows the notification data flow to the notification service. $S_{n}$ shows the $n^{\text{th}}$ scene where a suspicious object has been detected. The red line shows the notification data flow when detecting object anomalies.
Closed-Circuit Television (CCTV) cameras are widely applied in public spaces, traditionally serving as passive surveillance tools aimed primarily at security and monitoring (Ashby 2017). However, these systems often underutilized the potential of visual data they capture, missing opportunities to contribute to broader civic applications such as urban planning, public health, and resource management (Angelidou et al. 2018). By transforming CCTV networks into Smart Video Solutions (SVS), we can unlock new functionalities that go beyond mere surveillance to enhance community living and environments (Javed et al. 2022).
SVS leverage advancements in AI and machine learning to analyze video data in real-time, providing actionable insights that can improve city infrastructure and community services (Ardabili et al. 2023b). For example, the results can be used to inform urban planners about pedestrian flows for better traffic management (Rashvand et al. 2023), help public health officials monitor social distancing (Usmani et al. 2023), and assist in the efficient allocation of community resources (Dechouniotis et al. 2020). Such applications demonstrate the potential of SVS to transform video data into valuable information for civic improvements.
While traditional research has focused on the development of these systems in controlled environments (Pazho et al. 2023a), it is crucial to transition and evaluate SVS in real-world settings to address challenges such as latency, scalability, and privacy concerns (Ardabili et al. 2023a; Xu et al. 2018). Moving from lab to field ensures that these systems are robust and adaptable, capable of handling the complexities of dynamic urban environments and diverse community needs (Noghre et al. 2023).
The architecture of the SVS is illustrated in Figure 1, which represents the end-to-end design (Ardabili et al. 2023b). It begins with the environment where video data is captured by existing CCTV cameras. The data is then processed by the AI module, where computer vision techniques are applied for human behavior analysis. The processed information is sent to the server module for further data analysis and aggregation (Pazho et al. 2023a). To ensure privacy, metadata is stored in the cloud module, which also hosts a mobile application. This setup allows end-users to receive notifications and access insights on their devices, fostering real-time engagement and decision-making.
This article introduces the deployment and evaluation of an AI-enabled SVS designed to enhance safety and civic functionalities in community spaces such as educational areas, parking lots, and smart city applications. By employing privacy-preserving techniques, the system utilizes pose based data for human behavior analysis, aligning with ethical standards. The SVS applies innovative data visualization tools, including Occupancy Indicators, Statistical Anomaly Detection, Bird’s Eye View, and Heatmaps, to provide insights into pedestrian behaviors and resource allocation, thereby aiding urban planning and enhancing public safety.
The SVS has been implemented in a community college to evaluate its real-world effectiveness. Our findings demonstrate the system’s capability to manage 16 CCTV cameras with a throughput of 16.5 frames per second (FPS) over a 21-hour period, maintaining an average latency of 26.76 seconds for detecting behavioral anomalies and notifying users.
In summary, the contributions of this article are:
  • Adopting an AI-enabled SVS with existing CCTV infrastructure in real world, emphasizing civic applications beyond traditional security measures.
  • Presenting innovative data representation techniques that enhance urban planning and resource allocation by providing insights into pedestrian behaviors.
  • Providing a comprehensive real-world evaluation of the SVS, demonstrating its effectiveness, scalability, and privacy-preserving capabilities in community settings.
Related Works
Recent advancements in AI-driven video systems have leveraged state-of-the-art algorithms and system designs to perform high-level AI tasks using computer vision. Ancilia, for instance, employs a multi-stage computer vision pipeline aimed at enhancing public safety. While we have implemented this system as the baseline for our deployment, it is important to note that the original study reports the results in controlled environment. (Pazho et al. 2023a).
The deployment of SVS in real-world settings is essential to validate their effectiveness and address challenges such as latency, scalability, and privacy. While traditional research has focused on laboratory-based studies for real-time object tracking and anomaly detection (Pazho et al. 2023b; S. et al. 2021), there is a growing recognition of the need to evaluate these systems in practical environments. Real-world evaluation ensures that this technology can adapt to dynamic urban settings and meet diverse community needs.
Recent advancements in machine learning have enabled smart video technology to perform complex data analysis, offering insights for civic applications. For instance, Pramanik et al. demonstrated lightweight models for edge devices to reduce latency in urban environments (Pramanik, Sarkar, and Maiti 2021). Singh et al. introduced an AI-based video technology for fall detection, highlighting the potential for applications beyond traditional safety measures (Singh et al. 2023).
In 2023, the E2E-VSDL method by Gandapur utilized BiGRU and CNN for anomaly detection, achieving high accuracy (Gandapur 2022). Similarly, TAC-Net’s deep learning approach excelled in addressing anomaly scenarios (Huang et al. 2021).
Integrating these systems into smart city applications bridges the gap between theoretical research and practical implementation. Alshammari et al. developed an SVS using surveillance cameras in real-world settings (Alshammari and Rawat 2019), and RV College of Engineering Bengaluru improved object detector accuracy (Franklin and Dabbagol 2020). These efforts highlight the necessity of robust testbed support for evaluating SVS, addressing scalability and privacy challenges (Ma et al. 2019).
The combination of machine learning and IoT devices has transformed urban planning, providing insights for managing congestion and enhancing civic participation. Ardabili et al. introduced an end-to-end system design that integrates Ancilia with an end-user device for direct communication. We have adopted this comprehensive system design for our study’s deployment, allowing us to assess its effectiveness in practical, real-world environments (Ardabili et al. 2023b). Mahdavinejad et al. emphasized machine learning’s role in forecasting urban congestion (Mahdavinejad et al. 2018), while Zanella et al. advocated for accessible urban IoT data (Zanella et al. 2014). These studies illustrate how data-driven approaches contribute to civic applications.
Figure 2: End-to-end detailed system. $C_{0}$ represents the camera ID and for each camera, one AI Module $N_{0}$ including multi-AI-vision-models pipeline is assigned with. All the AI Modules send processed data to one Server Module database, and Server Module will re-identify the human track ID based on the feature extractor data. The statistical analyzer analyzes all the data in the database stored across all the cameras and communicates the results with Cloud Module. Cloud-native services are utilized to host the end users’ applications.
Figure 3: Locations of Cameras on the Campus. A maximum of 16 cameras are used
Despite progress in SVS research, comprehensive testbed support for evaluating real-world system performance remains limited. Our study addresses this gap by deploying a state-of-the-art SVS in a community college, demonstrating its capabilities and potential for enhancing civic applications in public settings.
Software System Features
The AI-based real-time video solution seamlessly integrates with existing CCTV infrastructures, creating a Physical-Cyber-Physical (PCP) system that delivers actionable information to end users. Such system consists of four key components: AI Modules, Server Module, Cloud Module, and End-user Devices, as illustrated in Figure 2. The design of the AI and server modules was inspired by work (Pazho et al. 2023a), while the cloud module and user device components were adapted from another study (Ardabili et al. 2023b).
AI Module/Modules
is a multi-stage pipeline optimized for real-time computer vision tasks. It processes image data in batches of 30 frames, where an object detector identifies objects and flags anomalies (Pazho et al. 2023a). A tracking algorithm then creates tracklets, and pose estimator extracts 2D skeletal data for individuals. Object anomaly detection and behavioral anomaly detection (Noghre, Pazho, and Tabkhi 2024; Pazho et al. 2024) are performed within this module respectively, with alerts sent to end-user devices for real-time response (Pazho et al. 2023b; Yao et al. 2024).
In our feature, YOLOv8 (Jocher, Chaurasia, and Qiu 2023), Bytetrack (Zhang et al. 2022), HRNet (Sun et al. 2019), GEPC (Markovitz et al. 2020a) and OSNet (Zhou et al. 2019) are used for further evaluation. The detailed decision-making process is provided in the supplementary materials for replication and further investigation.
Server Module
centralizes metadata and historical data from AI modules, stored in a MySQL database as outlined in Table 1. It handles global tracking and statistical analysis, using cosine similarity to re-identify individuals across cameras to analyze patterns (Pazho et al. 2023a). The design prioritizes privacy by avoiding raw data transmission to the Cloud Module.
Table 1: Example visualization of the database at Server Module
Record
Time
Camera Class ID
Bounding
Box
Feature Local ID Global ID
Anomaly
Score
00:00:00 1 0 [x, y, w, h] Tensors 15 1001 40
00:00:01 2 0 [x, y, w, h] Tensors 21 1001 40
23:59:59 1 0 [x, y, w, h] Tensors $9$ 1001 40
Cloud Module
leverages cloud-native services for data storage, management, and API generation (Dahunsi, Idogun, and Olawumi 2021). It minimizes lag between anomaly detection and notification, using a rule-based messaging service to send real-time alerts via email, text, or app notifications. A low-latency database supports real-time data access, while APIs generated through an application development kit optimize data retrieval (Mohammed et al. 2022; Ardabili et al. 2023b).
End User Devices
is designed to promptly notify users of detected anomalies via a smartphone application (Ardabili et al. 2023a). The app provides real-time data and analysis, ensuring consistent functionality across different devices and operating systems, thereby enhancing accessibility and delivering information efficiently.
Figure 4: This represents an example of 4 pipelines running in the testbed with IP cameras deployed at different locations.
Deployment and Setup
In lab settings, each advanced software features demonstrates significant potential individually; however, to fully realize its benefits, a comprehensive end-to-end evaluation in a real-world environment is essential. Such an evaluation allows for assessing the system’s load handling and endurance under actual conditions, ensuring that it can perform reliably and effectively when deployed, which might not be evident in a controlled environment.
In our study, an existing network of 16 AXIS IP cameras operating at 30 FPS with 720p resolution across the college campus is applied, covering approximately 35,000 square feet indoors and 60,000 square feet outdoors. Figure 3 visually represents the camera placements. Three cameras monitor outdoor areas, while the remaining 13 oversee various indoor locations, including entry points, vending machines, hallways, and communal spaces. The indoor cameras are mounted at 7 feet 6 inches, and the outdoor cameras at 10 feet 8 inches. Each camera has a varifocal lens with a horizontal field of view of 100 to 36 degrees. The system operates on a dedicated server equipped with a 16-core CPU, 252 GB of memory, and four GPUs, each with 24 GB VRAM.
As shown in Figure 4, a sample output from the Local Node during a scenario with four operational cameras is presented, with privacy-preserving instance segmentation applied.
Applications and Visualizations
This section illustrates how the data collected from real world can be analyzed and visualized to enhance situational awareness in applications such as urban planning, resource allocation, and crowd management.
Descriptive Data
In our descriptive analysis, the Global ID from human feature data serves as the unique identifier for tracking across data streams. We present five key metrics to provide a general overview of traffic flow and occupancy trends: real-time count of people, hourly average number of people per camera per location, total number of people and peak hour analysis over time. These metrics help to monitor occupancy, understand distribution trends, and identify peak traffic times, aiding in resource allocation and emergency planning.
Situational Awareness
Situational awareness is crucial for effective decision-making across various domains. Our study explores four following key visualization techniques:
Occupancy Indicator
provides insight into the number of people in a specific location by comparing current occupancy levels to historical data, which particularly useful in emergencies or public health scenarios. To calculate occupancy, the system compares current detections with historical percentiles, categorizing occupancy as ”Low,” ”Normal,” or ”High.”
Figure 5: Comparing different views of one camera. The graph (a) generated directly from database shows the original camera view. The graph (b) represents the average bird’s eye view coordinates.
(a) Camera 2- Weekday
(b) Camera 2- Weekend
(c) Camera 7- Weekday
(d) Camera 7- Weekend
Figure 6: Heatmap representation of the two cameras during weekday and weekends. Graph (a) and (c) represent the heatmaps during weekdays, while graph (b) and (d) show congestion of the same areas during weekends.
Statistical Anomaly
often indicate unusual traffic flow. By analyzing historical data, the system establishes a baseline, identifying deviations as anomalies. The system calculates such by comparing current detections with historical data, flagging events that exceed two standard deviations from the mean. In our analysis, zero detections were excluded to avoid skewing the data.
Bird’s Eye View
provides an accurate spatial representation, ideal for crowd management and area planning. The system normalizes object dimensions and uses a scale factor to accurately position objects within this view. Figure 5 compares the original and processed Bird’s Eye View from one camera, showing how this technique refines and clarifies spatial data.
Heat map
visually represent data distribution, useful for monitoring congestion and spatial usage. The system generates 2D heatmaps from Bird’s Eye View coordinates, applying Gaussian smoothing for clarity. Figure 6 presents heatmaps for different 24-hour intervals. Graph (a) and (c) show weekday pedestrian movement, while graph (b) and (d) represent the weekend congestions at the same areas. Comparing the density pattern during weekday and weekend, reflect differences in area usage.
The detailed algorithms for all visualization tools are available in the supplementary materials for replication and further investigation.
Community Engagement
To gauge public perception of such solution, we conducted seven rounds of survey studies in six public places across Charlotte, NC, during July and August 2023. A total of 410 respondents participated, providing insights into their concerns about safety and surveillance.
The results indicated that respondents were most concerned about safety in parking lots (54%), public transit (44%), and entertainment venues (42%). Additionally, significant concerns were raised about potential biases and discrimination in current passive surveillance technologies, with these concerns being important or very important to respondents across all demographic groups, highlighting a widespread sensitivity to this issue.  [^1] This study has been previously published, but in adherence to the double-blind review policy of the AAAI conference, the reference will be provided in the camera-ready version upon acceptance.
Interestingly, vulnerable groups, such as females and people of color, were statistically more likely to agree with the statement, ”The presence of video surveillance makes me feel safer.” Furthermore, respondents found our AI-driven video solution more beneficial, with over 50% rating it as very or extremely beneficial, compared to current passive technologies, which 42% rated as not effective or only moderately effective. Our solution also raised fewer privacy concerns, with 66% of respondents reporting no privacy concerns or only moderate concerns. These findings suggest a strong preference for AI-driven surveillance solutions over traditional passive technologies, especially in public places.
System Evaluation and Results
Figure 7: Latency and throughput trends concerning crowd densities during a week-long period for 8 camera nodes
(a) Latency Trends
(b) Throughput Trends
Figure 8: Latency and Throughput trends with respect to crowd densities across different nodes number running in parallelism.
Load Stress Evaluation
This section evaluated the system’s performance under increasing input loads by varying crowd densities from videos. Key metrics such as average latency and throughput were measured as we scaled the system using different numbers of pipeline nodes—one, four, eight, and twelve—to assess scalability and adaptability under increased workloads.
To rigorously assess the system’s performance under varying crowd densities and parallel node operations, we prepared ten videos with density levels ranging from 0 to 9 from CHAD dataset (Pazho et al. 2022). Each video lasted 150 seconds and operated at a frame rate of 30 frames per second. To ensure the statistical validity of our results, we averaged metrics over the central 100 batches in each run, excluding the initial 25 batches for system warm-up and the last 25 batches for cool-down effects.
Figure 8 graphically presents the trends for throughput and latency. The X-axis, labeled ”Density” (count), represents the average number of humans detected per batch (comprising 30 frames) for each input node number in the experiment. With one and four camera nodes, our system maintained latency under 10 seconds, showcasing stability regardless of scene crowdness. However, with eight nodes, latency increased beyond 20 seconds at higher density levels, particularly from level 5 onwards. A notable decrease in latency was observed with 12 nodes, especially at density levels 8 and 9, correlating with processing 108 individuals (12 nodes x 9 density).
Throughput consistently declined linearly with increasing density and node count, reaching a low of 4.56 FPS at the highest density level 9 with 12 nodes. The performance decrease is primarily due to the computational demands of pose estimator (Sun et al. 2019) processing high-density keypoints and is further constrained by CPU computation limits. As node count and density increase, these factors collectively create a bottleneck, leading to the observed decline in throughput and latency.
Real-World Endurance Evaluation
This evaluation assessed the system’s long-term stability using all 16 cameras in the testbed. We conducted extended trials with eight, twelve, and sixteen nodes running continuously for 21 hours and a week-long test using eight nodes to ensure the system’s robustness. This setup aimed to validate the system’s performance in realistic, extended operational conditions.
  • Eight-Node Setup: Covers key indoor areas, including entry points, hallways, and common spaces.
  • Twelve-Node Setup: Adds outdoor monitoring of parking lots.
  • Sixteen-Node Setup: Utilizes all cameras for comprehensive coverage.
As shown in Figure 7, the system’s performance over a week, running with eight cameras, demonstrated consistent latency and throughput results about crowd density. Specifically, the latency ranged from 2.5-7.8s, and throughput varied from 22.3-29.8 FPS. These findings confirm that irrespective of crowd density, the latency increased. At the same time, the FPS decreased, aligning with expectations and underscoring that the length of the videos does not significantly impact the system’s performance.
It is essential to note that the specific data spikes observed within this evaluation could stem from various factors, including network irregularities within the environment (as the input streams originate from IP cameras) or CPU memory usage nearing its limits. However, in light of the comprehensive experiment, these occasional data point spikes remain within acceptable bounds. The system consistently maintained optimal performance levels, showcasing its robustness and capacity to effectively manage and adapt to various operational scenarios.
For configurations with 8, 12, and 16 camera nodes, we observed similar trends over the 21-hour-long experiment with varying performance metrics: average latencies ranged from 2.6-4.8s, 5.3-6.5s, and 6.7-10.5s, respectively, while corresponding throughputs ranged from 28.5-26.5 FPS, 20.5-18 FPS, and 16.5-14.5 FPS. Detailed results with visualizations for 8, 12, and 16 camera nodes are provided in the supplementary materials for replication and further investigation.
(a) Object Anomaly
(b) Behavior Anomaly
Figure 9: Represent latency rug plot with different anomaly scenarios. 50 data points for each camera number
Physical-Cyber-Physical Evaluation (Anomaly detection)
This section evaluates the system’s Physical-Cyber-Physical (PCP) latency, representing the end-to-end time from when an anomaly appears to when the end user receives a notification. We assessed both object and behavioral anomaly detection (Markovitz et al. 2020b) using four, eight, twelve, and sixteen AI pipeline nodes reading input frames from the same camera node, conducting twenty experiments per anomaly type and collecting 400 data samples. The latency measures the time from detecting anomalies, such as high-priority objects or dangerous behaviors like fighting, to alerting the end user. A three-second transmission delay from cameras to AI Nodes was observed, slightly higher than other IP systems, due to our unique network setup. This delay is included in the overall PCP measurements, though manual recording of latency data may introduce minor discrepancies.
The results indicate that increasing the number of cameras significantly impacts PCP latency. For example, object anomaly latency increased from 4.7 seconds with four cameras to 19.68 seconds with sixteen, while behavioral anomaly latency rose from 9.34 seconds to 26.76 seconds. This increase is attributed to system traffic, multitasking overhead, and the processing capacity of the system.
Figure 9 displays the distribution of 400 data points from our PCP experiments. Graph (a) shows the distribution for object anomaly detection, with less variation in the four-camera setup compared to the sixteen-camera setup. Graph (b) illustrates the distribution for behavioral anomaly detection, where the data is more scattered with an increased number of cameras. The standard deviation for object detection latency increased from 0.65 to 1.12 seconds as the camera count increased, indicating a slight broadening of latency due to system complexity and resource contention.
This evaluation demonstrates the system’s capability to enhance safety and operational efficiency in real-world scenarios. The ability to notify end users within approximately 20 seconds is crucial for managing anomalies in public spaces, aiding in the prevention of potential dangers and improving resource allocation.
Overall, these real-world evaluations allows research how these key parameters, such as the number of cameras and crowd density, could affect system performance, such as latency and throughput in different public safety scenarios in case of emergency, is crucial in decision making and taking the proper actions. For example, retail and transport hubs could analyze foot traffic and crowd density, offering insights into peak hours and emergency response planning (Hosseini et al. 2024).
Conclusion
This paper presented a comprehensive evaluation of an AI-based real-time video solution, integrated with existing CCTV infrastructures, to enhance situational awareness through object and behavioral anomaly detection. The system was tested across various configurations, including up to sixteen camera nodes, to assess its performance under real-world conditions.
Our evaluation highlighted the system’s ability to maintain operational efficiency with increasing node counts, despite a corresponding rise in PCP latency. Object and behavioral anomaly detection latencies ranged from 4.7 seconds with four cameras to 26.76 seconds with sixteen, reflecting the system’s scalability and robustness. Through extensive testing, the system demonstrated its capability to deliver timely notifications to end users, crucial for managing anomalies in public spaces. This work underscores the importance of real-world evaluations in optimizing AI-driven video solutions, ensuring they meet the demands of dynamic environments and contribute to enhanced public safety and operational efficiency.